### v2

# Scalable Bayesian Image-on-Scalar Regression for Population-Scale Neuroimaging Data Analysis

Abstract
Bayesian Image-on-Scalar Regression (ISR) offers significant advantages for neuroimaging data analysis, including flexibility and the ability to quantify uncertainty. However, its application to large-scale imaging datasets, such as found in the UK Biobank, is hindered by the computational demands of traditional posterior computation methods, as well as the challenge of individual-specific brain masks that deviate from the common mask typically used in standard ISR approaches. To address these challenges, we introduce a novel Bayesian ISR model that is scalable and accommodates inconsistent brain masks across subjects in large-scale imaging studies. Our model leverages Gaussian process priors and integrates salience area indicators to facilitate ISR. We develop a cutting-edge scalable posterior computation algorithm that employs stochastic gradient Langevin dynamics coupled with memory mapping techniques, ensuring that computation time scales linearly with subsample size and memory usage is constrained only by the batch size. Our approach uniquely enables direct spatial posterior inferences on brain activation regions. The efficacy of our method is demonstrated through simulations and analysis of the UK Biobank task fMRI data, encompassing 38,639 subjects and over 120,000 voxels per image, showing that it can achieve a speed increase of 4 to 11 times and enhance statistical power by 8% to 18% compared to traditional Gibbs sampling with zero-imputation in various simulation scenarios.
Keywords: Image-on-Scalar regression; Scalable algorithm; Memory-mapping; UK Biobank data; Individual-specific masks.
1 Introduction
Magnetic Resonance Imaging (MRI) is a non-invasive technique renowned for its comprehensive insight into the brain’s structure and function. Functional MRI (fMRI) is an imaging modality that detects neuronal activity via fluctuations in the blood oxygen level dependent (BOLD) signal, providing insights into brain activity (Lindquist, 2008). Task-based fMRI can be used to identify regions where individual traits (e.g. cognitive ability) associate with brain function. We aim to map the influence of a single trait on activity over the whole brain. This represents a classical problem in imaging statistics, where the outcome is an image, and the predictors are multiple scalar variables, commonly known as Image-on-Scalar regression (ISR). As an example, using UK Biobank data (Sudlow et al., 2015), we focus on examining the influence of age on brain activity across the whole brain using task fMRI data. In this scenario, the task fMRI data is the outcome image, while age is the scalar predictor variable.
The analysis of large-scale brain fMRI data presents significant challenges including low signal-to-noise ratio and complex anatomical brain structure (Lindquist, 2008; Smith and Nichols, 2018). The advent of large-scale neuroimaging studies, such as the UK Biobank and Adolescent Brain Cognitive Development (ABCD) study, has introduced new computational challenges for traditional statistical tools in analyzing large-scale fMRI data. These challenges arise from the substantial size of the datasets, as well as the scalability of posterior computation algorithms and the difficulties encountered in achieving convergence in high-dimensional settings.
A specific challenge associated with the UK Biobank fMRI data is the presence of individual-specific brain masks. The brain typically occupies less than half of the cuboid image volume, and a brain mask is used to identify which voxels constitute brain parenchyma and should be included in the analysis. Even after registration of brain images to standard space, each subject’s fMRI data can have a unique brain mask, and this is the case in the UK Biobank task fMRI data.
In this paper, we seek to address these challenges by presenting a Bayesian hierarchical model for Image-on-Scalar regression (ISR) that incorporates a sparse and spatially correlated prior on the exposure coefficient. Furthermore, we propose an efficient algorithm utilizing Stochastic Gradient Langevin Dynamics (SGLD)(Welling and Teh, 2011) to ensure scalability. To accommodate the varying individual-specific masks, we employ imputation techniques, enabling us to analyze a wide spatial mask across all individuals.
1.1 UK Biobank Data
The use of imaging biomarkers in clinical diagnostics and disease prognostics has been historically hindered by the lack of imaging data collected before disease onset. The UK Biobank is collecting longitudinal data on one million UK residents of which a sub-sample of one hundred thousand are being imaged longitudinally (Miller et al., 2016). The UK Biobank data provides multimodal brain imaging data including structural, diffusion, and functional MRI data. We focus on the task fMRI data, an emotion task where participants are asked to identify faces with negative emotions using shape identification as the baseline task. The objective of this emotional task is to actively involve cognitive functions ranging from sensory and motor areas to regions responsible for processing emotions. Recent studies using the UK Biobank have utilized multimodal data to predict brain age (Cole, 2020), have analyzed the association between resting state connectivity data, education level, and household income (Shen et al., 2018), and have applied deep learning to sex classification in resting state and task fMRI connectomes (Leming and Suckling, 2021); amongst others. (Elliott et al., 2018; Littlejohns et al., 2020).
1.2 Traditional and Recent Practices in ISR
The most common practice for Image-on-Scalar regression is the mass univariate analysis approach (MUA) (Groppe et al., 2011). Although MUA is computationally efficient, easy to implement, and has well-developed multiple comparison correction methods to control false discovery rates, MUA ignores spatial structure and tends to have low statistical power when the data has a low signal-to-noise ratio. To account for spatial dependence, as discussed in Morris (2015), a common approach is to use a low-rank approximation. Built on the principle components idea to utilize spatial correlation, Ramsay and Silverman (2005) and Reiss et al. (2010) proposed a penalized regression method using basis expansion to reduce the dimension of the functional outcome. Zhu et al. (2014) propose another model that uses local polynomial and kernel methods for estimating the spatially varying coefficients with discontinuity jumps. Yu et al. (2021) and Li et al. (2021) use bivariate spline functions to estimate the spatial functional estimator supported on the 2-dimensional space. Zhang et al. (2022) propose a quantile regression method that can be applied to high-dimensional outcomes. A common issue with all the aforementioned frequentist methods is that it can be difficult to make inferences on the active area selection based on these penalized low-rank models. Zhang et al. (2023) develop an efficient deep neural network approach to estimate the spatially varying parameters of very high-dimension with complex structures, but they only focus on point estimation rather than statistical inference. Zeng et al. (2022) propose a Bayesian model with a prior composed of a latent Gaussian variable and a binary selection variable to account for both sparsity and spatial correlation. However, the dense covariance matrix can require large memory and computational power when the outcome is very high-dimensional. To address these limitations, we propose a scalable Bayesian Image-on-Scalar regression model where the functional coefficient is assigned a sparse and spatially correlated prior so that we can make inferences on the activation areas directly from the posterior inclusion probability (PIP).
1.3 Subject-specific Masks in Brain Imaging
The brain mask used for a multisubject fMRI analysis is typically the intersection of the individual-specific masks, only analyzing voxels where all subjects have data. However, in the UK Biobank, an intersection mask of 38,639 subjects reduces the analysis volume by 71% relative to the average subject mask volume. Some authors will attempt to minimize this effect by identifying subjects with particularly small masks, but this is a laborious process and discards data. For a particular subject, all voxels that lie in the set difference of the union mask and that subject’s mask results in missing values for said subject. Many of the voxels with missing data occur around the edge of the union mask (Mulugeta et al., 2017). Historically, researchers usually do not account for missing data, however, for PET data, missing values are imputed using a soft mean (Hammers et al., 2007) derived using available data at each voxel.
In volumetric fMRI data, individual brain masks can be slightly different from one another, due to some subject’s brain falling outside of the field-of-view truncation, residual variation in inter-subject brain shape not accounted for by atlas registration, and susceptibility-induced signal loss. In particular, voxels above the nasal cavity and above the ear canals suffer from signal loss and are classified as non-brain tissue in a subject-specific manner.
All brain image analyses requires a brain mask to avoid wasted computation and uninterpretable results on non-brain voxels. For mass-univariate fMRI analyses, some authors created custom software (Szaflarski et al., 2006; Maullin-Sapey and Nichols, 2022) or used mixed effect models with longitudinal fMRI data on each voxel (Szaflarski et al., 2012) to account for subject-specific masks, or explicitly accounted for missing data with multiple imputation (Vaden Jr et al., 2012). Among the major fMRI software packages, however, neither FSL nor SPM can account for subject-specific masks, and only AFNI’s 3dMEMA (Chen et al., 2012) (simple group analysis) or 3dLME (Chen et al., 2013) (general mixed effects) accounts for subject-specific masks. However, all of these methods are mass-univariate. To the best of our knowledge, our work is the first spatial Bayesian method to account for varying missing data patterns over the brain. By using individual masks with imputation, we make the most use of all collected data.
1.4 Scalable Posterior Algorithms
The scalable posterior algorithm we use is based on the stochastic gradient Langevin dynamics (SGLD) algorithm (Welling and Teh, 2011). The SGLD algorithm is effective at handling large-scale data as it approximates the posterior gradient using subsamples of the data. Variations on the SGLD algorithm for scalable posterior sampling have been proposed. Wu et al. (2022) proposed a Metropolis-Hasting algorithm using mini-batches where the proposal and acceptance probability are both approximated by the current mini-batch. Kim et al. (2022) proposed an adaptive SGLD algorithm (Adam SGLD) that sets a preconditioner for SGLD and allows the gradient at different directions to update with different step sizes. Aside from these MCMC algorithms, variational Bayesian inference (Jaakkola and Jordan, 1999) is another popular option for approximating the posterior mean in high-dimensional settings. There has been increasing use of variational inference for high-dimensional posteriors such as imaging data analysis (Kaden et al., 2008; Kulkarni et al., 2022), and various scalable extensions of variational inference (Hoffman et al., 2013; Ranganath et al., 2014; Blaiotta et al., 2016). Although these variational methods are computationally efficient, our goal, however, is to obtain the entire MCMC sample that provides uncertainty quantification for the activation areas of interest. Hence, we settled on the SGLD-type algorithm for its scalability.
The main contributions of our proposed method are
  1. 1.
    to provide an efficient posterior computation algorithm for Bayesian Image-on-Scalar regression, scalable to large sample size and high-resolution image;
  2. 2.
    to introduce the individual-specific brain masks and expand the analysis region from an intersection mask of all individuals to an inclusive mask using imputation.
In particular, our method uses batch updates and memory-mapping techniques to analyze large sample imaging data that is too big to fit into random access memory on many computers. In addition, we provide an imputation-based method that allows us to handle individual-specific masks and makes full use of the observed data.
The rest of this paper is organized as follows. In Section 2, we introduce our Bayesian Image-on-Scalar model. In Section 3, we discuss algorithm and computational details. In Section 4, we demonstrate the performance of our proposed algorithms against the classical MUA approach under different simulation settings. In Section 5, we apply the proposed method to the currently available subset of the UK Biobank imaging data and provide sensitivity analysis results. In Section 6 we summarize our contribution and provide a general discussion. Additional simulation and real data results can be found in the Supplementary Materials. Our implementation is provided as an R package, SBIOS, and is publicly available on Github 111 See the SBIOS R package on Github page https://github.com/yuliangxu/SBIOS or in the Supplementary.
2 Model
Let $\mathcal{B}\subset\mathbb{R}^{3}$ denote the entire brain region. Let $\left\{s_{j}\right\}_{j=1}^{p}\subset\mathcal{B}$ be a set of fixed grid points in $\mathcal{B}$, on which we observe brain image intensity values. For individual $i~{}(i=1,\dots,n)$, let $Y_{i}(s_{j})$ be the image intensity at voxel $s_{j}$. To incorporate the individual-specific masks, let $\mathcal{V}_{i}$ denote the set of locations where the image intensity is observed for individual $i$, i.e., for any $s_{j}\in\mathcal{V}_{i}$, $Y_{i}(s_{j})$ is not missing. For any $i=1,\dots,n$, $\mathcal{V}_{i}\subset\left\{s_{1},\dots,s_{p}\right\}.$ Let $X_{i}$ be the primary covariate of interest, $Z_{ik}$ be the $k$-th confounding covariate for $k=1,\dots,q$. We propose an Image-on-Scalar regression model. For individual $i$ and any $s_{j}\in\mathcal{V}_{i}$,
$Y_{i}(s_{j})=X_{i}\beta(s_{j})\delta(s_{j})+\sum_{k=1}^{m}\gamma_{k}(s_{j})Z_{ik}+\eta_{i}(s_{j})+\epsilon_{i}(s_{j}),\quad\epsilon_{i}(s_{j})\sim\mathrm{N}(0,\sigma_{Y}^{2}).$ (1)
The spatially varying parameter $\beta(s)$ estimates the magnitude in the image intensity that can be explained by the predictor $X$, and the binary selection indicator $\delta(s)$ follows a Bernoulli prior with selection probability $p(s)$. In practice, we set $p(s)=0.5$ for any $s\in\mathcal{B}$ as the prior for $\delta(s)$. The selection variable $\delta(s)$ determines the active voxels in the brain associated with the predictor $X$ and $\beta(s)\delta(s)$ is of main interest. The spatially varying parameter $\gamma_{k}(s)$ is the coefficient for the $k$-th confounder $Z_{k}$, and $\eta_{i}(s)$ accounts for individual level spatially correlated noise. By introducing the individual effect $\eta_{i}$ as a parameter, we are separating spatially correlated noise from spatially independent noise $\epsilon_{i}$ and we can safely assume a completely independent noise term $\epsilon_{i}(s)$ across all locations $s_{j}$, hence avoiding large-scale covariance matrix computations in the noise term. This is similar to the correlated noise model in Zhu et al. (2014). Proposition 1 in Zhang et al. (2023) provides sufficient conditions for model identifiability of model (1): 1) the design matrix $\tilde{\mathbf{X}}:=\left(\mathbf{X},\mathbf{Z}\right)\in\mathbb{R}^{n\times(m+1)}$ is a full rank matrix, and 2) for any $i$ and any $s\in\mathcal{B}$, denote $\boldsymbol{\eta}(s)=\left(\eta_{1}(s),\dots,\eta_{n}(s)\right)\in\mathbb{R}^{n}$, $\tilde{\mathbf{X}}^{T}\boldsymbol{\eta}(s)=0$.
For model (1), we specify the following priors:
$\delta(s)$ $\sim\mathrm{Ber}\{p(s)\},\quad\mbox{ for any }s\in\mathcal{B}$ (2)
$\beta(s)$ $\sim\mathcal{GP}(0,\sigma_{\beta}^{2}\kappa),\quad\mbox{ for any }s\in\mathcal{B}$ (3)
$\gamma_{k}(s)$ $\sim\mathcal{GP}(0,\sigma_{\gamma}^{2}\kappa),\quad\mbox{ for any}s\in\mathcal{B},\quad j=1,\dots,q$ (4)
$\eta_{i}(s)$ $\sim\mathcal{GP}(0,\sigma_{\eta}^{2}\kappa),\quad\mbox{ for any }s\in\mathcal{B}$ (5)
The spatially-varying functional coefficients $\beta(s),\gamma_{k}(s),\eta_{i}(s)$ are assumed to have Gaussian Process (GP) priors with mean 0 and kernel function $\sigma^{2}\kappa(\cdot,\cdot)$, where $\sigma^{2}$ can be different for each functional parameter. Popular choices of the kernel function $\kappa$ including the exponential square kernel and the Matérn kernel (8). We use the latter in both simulation studies and real data analysis as it offers flexible choices of the kernel parameters. Here, all GPs are assumed to have the same kernel function $\kappa$ for computational efficiency. In addition, we expect the individual effects $\eta_{i}$ to be smoother than $\beta$, hence choosing an appropriate kernel for $\beta$ is sufficient for the estimation of $\eta_{i}$. In the UK Biobank data analysis, $\kappa$ is chosen to reflect the outcome image $Y_{i}(s)$ correlation structure in a data-adaptive way, as discussed in Section 5.
3 Posterior Computation
3.1 Posterior Sampling with Gaussian Process Priors
The 3D task fMRI image data is divided into $R$ regions using the Harvard-Oxford cortical and subcortical structural atlases (Desikan et al., 2006). Between-region independence is assumed when constructing the Gaussian kernel for $\beta$, $\gamma_{k}$ and $\eta_{i}$. Instead of assuming a whole brain correlation structure for the GP priors, a block diagonal covariance structure is computationally efficient and allows us to capture more detailed information within each region, especially regions of smaller size and complex spatial structures.
To sample from the posterior of the GPs, we use a basis decomposition approach. By Mercer’s theorem (Rasmussen and Williams, 2005), for any $\beta(s)\sim\mathcal{GP}(0,\sigma_{\beta}^{2}\kappa)$, we can use a basis decomposition,
$\beta(s)=\sum_{l=1}^{\infty}\theta_{\beta,l}\psi_{l}(s),\quad\theta_{\beta,l}\sim\mathcal{N}(0,\sigma_{\beta}^{2}\lambda_{l}),$
[…]
With the basis expansion coefficient, we can sample from the $L$-dimensional space to model the $p$-dimensional image data. We also partition the brain into regions to further speed up computation and assume a region-independence structure for the spatially varying parameters $\beta,\gamma_{k}$, and $\eta_{i}$. Assume there are $r=1,\dots,R$ regions that form a partition of the mask $\mathcal{B}$, denoted as $\mathcal{B}_{1},\dots,\mathcal{B}_{R}$. For the three GP priors (3) - (5), we assume that the kernel function $\kappa(s_{j},s_{k})=0$ for any $s_{j}\in\mathcal{B}_{r},s_{k}\in\mathcal{B}_{r^{\prime}},r\neq r^{\prime}$, and the prior covariance matrix on the fixed grid has a block diagonal structure.
For the $r$-th region, let $p_{r}$ be the number of voxels in $\mathcal{B}_{r}$ and let $Q_{r}=\left(\psi_{l}(s_{r,j})\right)_{l=1,j=1}^{L_{r},p_{r}}\in\mathbb{R}^{L_{r}\times p_{r}}$ be the matrix with the $(l,j)$-th component $\psi_{l}(s_{r,j})$ where $\left\{s_{r,j}\right\}_{j=1}^{p_{r}}$ forms the fixed grid in $\mathcal{B}_{r}$. With the region partition, the GP priors on the $r$-th region can be reexpressed as $\beta_{r}=\left(\beta(s_{r,1}),\dots,\beta(s_{r,p_{r}})\right)^{T}\approx Q_{r}\boldsymbol{\theta}_{\beta,r}$, where $\boldsymbol{\theta}_{\beta,r}\sim\mathcal{N}\left(0,\sigma_{\beta}^{2}D_{r}\right)$ and $D_{r}$ is a diagonal matrix with diagonal $(\lambda_{r,1},\dots,\lambda_{r,L_{r}})\in\mathbb{R}^{L_{r}}$.
To present the working model with region partitions, denote $Y^{*}_{i,r}=Q_{r}^{T}Y_{i,r}\in\mathbb{R}^{L_{r}}$ as the low-dimensional mapping of the $i$th image on the $r$-th region where $Y_{i,r}=\left\{Y_{i}(s_{j})\right\}_{s_{j}\in\mathcal{B}_{r}}\in\mathbb{R}^{p_{r}}$, and $\epsilon^{*}_{i,r}=Q_{r}^{T}\epsilon_{i,r}$, $\epsilon_{i,r}=\left\{\epsilon_{i}(s_{j})\right\}_{s_{j}\in\mathcal{B}_{r}}\in\mathbb{R}^{p_{r}}$. Let $\mathrm{diag}\left\{x\right\}$ be the diagonal matrix with diagonal $x$. Let $\boldsymbol{\delta}_{r}=\left\{\delta(s_{j})\right\}_{s_{j}\in\mathcal{B}_{r}}\in\mathbb{R}^{p_{r}}$. After basis decomposition
$Y_{i,r}^{*}=Q_{r}^{T}X_{i}\mathrm{diag}\{\boldsymbol{\delta}_{r}\}Q_{r}\boldsymbol{\theta}_{\beta,r}+\sum_{k=1}^{m}\theta_{\gamma,j,r}Z_{j}+\theta_{\eta,i,r}+\epsilon^{*}_{i,r}$ (6)
with the prior specification $\boldsymbol{\theta}_{\beta,r}\sim N(0,\sigma_{\beta}^{2}D_{r})$, $\boldsymbol{\theta}_{\gamma,k,r}\sim N(0,\sigma_{\gamma}^{2}D_{r})$, $\boldsymbol{\theta}_{\eta,i,r}\sim N(0,\sigma_{\eta}^{2}D_{r})$, and the low-dimensional noise $\epsilon^{*}_{i,r}\sim N(0,\sigma_{Y}^{2}I_{L,r})$. The working model (6) based on regionally independent kernels performs a whole brain analysis since $\sigma_{\beta}^{2},\sigma_{\gamma}^{2},\sigma_{\eta}^{2}$, and $\sigma_{Y}^{2}$ are estimated globally across regions. This is also the first step towards reducing memory cost by using a low-dimensional approximation. The finite cutoff $L$ is chosen to reflect the flexibility of the true functional parameters: fewer bases are required to approximate the smooth function $\beta(s)$. In our real data analysis, $L$ is determined by extracting the eigenvalues of the covariance kernel matrix. For one brain region, first compute the $p\times p$ dimensional covariance matrix with appropriately tuned covariance parameters, get the eigenvalues of such covariance matrix, and choose the cutoff such that the summation $\sum_{l=1}^{L}\lambda_{l}$ is over $90\%$ of $\sum_{l=1}^{p}\lambda_{l}$. Details for choosing the covariance parameters in (8) and sensitivity analysis on the $90\%$ cutoff are discussed in Sections 5 and 5.4.
3.2 Scalable Algorithm for Large Dataset
Inspired by Welling and Teh (2011) and Wu et al. (2022), we propose Algorithm 1 based on SGLD to sample the posteriors from a large data set. The SGLD algorithm outperforms Gibbs sampling (GS) in three ways: 1) under the assumption that all individuals are independently distributed according to the proposed model, the SGLD algorithm allows us to compute the log-likelihood on a small subset of data, making it computational efficient compared to GS; 2) the SGLD algorithm adds a small amount of noise to the gradient of the log posterior density, making it more effective in exploring the posterior parameter space; and 3) as the step size decreases to 0, the SGLD algorithm provides a smooth transition from the stochastic optimization stage to the posterior sampling stage.
We apply the SGLD algorithm on $\theta_{\beta}$. Denote $\theta_{\beta}^{(t)}$ as the value at the $t$-th iteration. Let $\tau_{t}$ be the step size at the $t$-th iteration, let $\pi(\theta_{\beta,l})$ denote the prior, and let $\pi_{i\in\mathcal{I}}\left(Y_{i}\mid X_{i},Z_{i},\theta\right):=\prod_{i\in\mathcal{I}}\pi\left(Y_{i}\mid X_{i},Z_{i},\theta\right)$ be the likelihood for a subsample $\mathcal{I}$, where $\theta$ is the collection of all parameters. Denote by $n_{s}$ the number of subjects in the subsample $\mathcal{I}$. Let $\nabla f(x)$ be the gradient of $f(x)$. At the $t$-th iteration, the $l$-th component in $\theta_{\beta}$ is updated as
$\theta_{\beta,l}^{(t)}\longleftarrow\theta_{\beta,l}^{(t-1)}+\frac{\tau_{t}}{2}\nabla\log\pi(\theta_{\beta,l}^{(t-1)})+\frac{\tau_{t}}{2}\frac{n}{n_{s}}\nabla\log\pi_{i\in\mathcal{I}}\left(Y_{i}\mid X_{i},Z_{i},\theta^{(t-1)}\right)+\sqrt{\tau_{t}}\varepsilon_{l}$ (7)
where $\varepsilon_{l}\overset{\mathrm{iid}}{\sim}N(0,1)$. Let $L_{r}$ be the number of basis coefficients for the $r$-th brain region. The time complexity for (7) is $\min\left\{O(L_{r}^{3}),O(L_{r}^{2}n_{s})\right\}$. The full sample size $n$ is usually significantly larger than $L_{r}$, hence approximating the full likelihood with a subsample of the data significantly decreases the computational complexity.
Based on the mini-batch idea of the SGLD algorithm, we propose the following Algorithm 1, where the large data set is first split into $B$ smaller batches. Each batch of data is loaded using memory-mapping techniques. Within each batch, a small subsample of size $n_{s}$ is randomly drawn to be used at each iteration. By splitting the full data into $B$ batches, we reduce the auxiliary space complexity for computing the required summary statistics down to the size of $B$, instead of the size of the full data.
Algorithm 1 Scalable Bayesian Image-on-Scalar (SBIOS) regression with memory mapping
1:Set subsample size $s$, an integer $t_{I}$ for the frequency to update $\eta_{i}$, and initial values for all parameters.
2:Split the entire sample into $B$ batches, sequentially load each batch of data and save each batch to the disk using memory-mapping. Set batch index $b$ to 1.
3:for  iteration $t=1,2,\ldots,T$ do
4:     Update the step size $\tau_{t}$ for $t$-th iteration.
5:     Load batch $b$ into memory.
6:     for  region $r=1,...,R$ do
7:         Randomly select a subsample $\mathcal{I}$ of size $n_{s}$ from batch $b$.
8:         Update $\theta_{\beta,l}$ for region $r$ using (7) based on the selected subsample.
9:     end for
10:     Update $\gamma,\delta,\sigma_{\gamma},\sigma_{\beta}$ using Gibbs sampling.
11:     $b=b+1$. If $b>B$, set $b=1$.
12:     if $t$ is a multiple of $t_{I}$ then
13:         Iterate through all batches to update $\{\eta_{i}\}_{i=1}^{N}$.
14:         Update $\sigma_{Y},\sigma_{\eta}$ using Gibbs sampling.
15:         (Optional) Impute the missing outcome $Y_{i}(s_{j})$ for all the missing voxel indices $s_{j}\notin\mathcal{V}_{i}$.
16:     end if
17:end for
On line 10 of Algorithm 1, using Gibbs sampling to update $\gamma,\delta$ also requires the entire data set, but the posterior distributions of $\gamma$ and $\delta$, in fact, only rely on summary statistics that can be pre-computed based on the entire data set. In practice, at the beginning of the algorithm, we iterate through each batch once to compute these summary statistics and directly use them to update $\gamma$ and $\delta$ at each iteration. The same cannot be done for $\theta_{\beta,l}$, because $\delta(s)$ can be 0 or 1, and the posterior variance for $\left(\theta_{\beta,1},\dots,\theta_{\beta,L}\right)^{T}$ depends on $Q^{T}\mathrm{diag}\left\{\delta\right\}Q$, which is no longer a diagonal matrix and requires updating at each iteration. See Equation (6). Hence, sampling $\theta_{\beta,l}$ is more computationally demanding than sampling $\theta_{\gamma,l}$. We provide detailed derivations of the posterior distribution for $\theta_{\beta,l},\theta_{\gamma,l},\delta(s)$ in the Supplementary Material, Section 1.
On line 13 of Algorithm 1, since $\theta_{\eta,i,l}$ is the coefficient for individual-level effect, we must iterate through all samples to update all of the $\theta_{\eta,i,l}$. As such, storing $\theta_{\eta,i,l}$ in memory grows as the sample size $n$ increases. As $\eta_{i}(s)$ is more of a nuisance parameter and not our main focus, we choose to update $\eta_{i}$ less frequently. Further improvements to memory allocation can be implemented if we also use batch-splitting on $\eta$ and save the samples of $\eta$ as file-backed matrices. This is implemented for the SBIOS0 function in our package as an example.
For the optional imputation found on line 15 of Algorithm 1, imputing $Y_{i}(s_{j})$ where $s_{j}\notin\mathcal{V}_{i}$ means that the missing values in $Y_{i}$ and all summary statistics associated with $Y_{i}$ need to be updated every $t_{I}$ iterations. Hence we use an index-based updating scheme. For each $i$, denote the complementary set $\mathcal{V}_{i}^{c}:=\left\{s_{j}\right\}_{j=1}^{p}-\mathcal{V}_{i}$ as the index set of all missing voxels for individual $i$. We keep track of $\mathcal{V}_{i}^{c}$, and create a vector Y_imp to store the imputed outcome values only on those indices in $\mathcal{V}_{i}^{c}$ for each $i$, and update the corresponding summary statistics every time $Y_{i}(s_{j}),s_{j}\in\mathcal{V}_{i}^{c}$ is updated. The time complexity to update all missing values in $Y_{i}$ is $O(L\times(|\mathcal{V}_{i}^{c}|))$ where $|\mathcal{V}_{i}^{c}|$ is the number of missing voxels for individual $i$, and $L$ is the total number of basis coefficients. We provide a detailed algorithm for updating missing outcomes in the Supplementary Materials.
Algorithm 1 is implemented using the Rcpp package bigmemory (Kane et al., 2013). The bigmemory package allows us to store large matrices on disk as a big.matrix class and extract the address of the large matrices. At the beginning of Algorithm 1, when the entire data is split into smaller batches, each batch is loaded in R as a big.matrix class, then the address of these big matrices is passed to Algorithm 1, accessing different batches of data becomes very efficient and memory-conserving.
4 Simulations
4.1 Simulation Design
We compare our proposed method with existing methods via simulation studies. We consider the following four different methods.
Mass Univariate Analysis (MUA) is one of the most commonly used methods for ISR models. MUA ignores any spatial correlation in the image and treats the data at all voxels as completely independent and performs linear regression on the data at each voxel. Ignoring spatial correlation may lead to potential overfitting and low power. When comparing selection accuracy for $\beta(s)\delta(s)$, we use the Benjamini-Hochberg adjusted p-values (Benjamini and Hochberg, 1995) to control the false discovery rate. The missing voxels are directly imputed as 0 in the MUA.
Bayesian Image-on-Scalar (BSIOS) model is the baseline Bayesian method for our proposed model (1)-(5), where all parameters are sampled via Gibbs sampling. Missing voxels are imputed as 0. The sampling algorithm reads the entire data set into memory.
Scalable Bayesian Image-on-Scalar model with 0 imputation (SBIOS0) is our proposed model (1)-(5) with sampling performed via the SGLD algorithm and memory-mapping. Missing data at voxels in individual masks are replaced by 0.
Scalable Bayesian Image-on-Scalar model with GS-imputation (SBIOSimp) is the proposed model (1)-(5) using the SBIOS0 algorithm, however, missing data in the individual masks are imputed from the proposed model.
(a) True covariate $\beta(s)\delta(s)$.
(b) Missing Pattern I
(c) Missing Pattern II
Figure 1: True signal and missing patterns. Missing Pattern I & II: OP ranges from 0.5 to 1.
To demonstrate the scalability of the proposed method, three simulation studies are used to show the selection accuracy (Simulation I), maximum memory usage (Simulation II), and time scalability (Simulation III). In each simulation, the entire sample is split into $B$ batches of data, each containing 500 subjects, with subsample size $n_{s}$ fixed at 200. The true image $\beta(s)$ is a 2D image of size $p=60\times 60=8100$, as shown in Figure 1 (a). The binary variable $\delta(s)$ is generated as 1 when $\beta(s)$ is nonzero, and 0 otherwise. The confounder parameters $\gamma_{k}(s)$ and individual effects $\eta_{i}(s)$ are generated based on the coefficients $\theta_{\gamma,k,l},\theta_{\eta,i,l}$, sampled from the standard normal distribution. Figure 1(b) and 1(c) are visual illustrations of the spatial observed proportion over all samples, where the image value at each pixel $s_{j}$ is $n^{-1}\sum_{i=1}^{n}I(s_{j}\in\mathcal{V}_{i})$. The difference between these 2 missing patterns is that missing pattern I selects part of the active area where $\beta(s)\neq 0$ with missing proportion of 50% while missing pattern II only selects missing voxels outside of the active area.
When generating the covariance kernel function, the $90\times 90$ 2D images are evenly split into 9 ($3\times 3$) regions that are assumed to be independent from each other, a priori. Hence we compute the basis decomposition of each region supported on a square of $30\times 30$ pixels, where the kernel function is the Matérn kernel with $\rho=2,\nu=1/5$ as in (8).
$\kappa(s^{\prime},s;\nu,\rho)=C_{\nu}(\|s^{\prime}-s\|_{2}^{2}/\rho),~{}C_{\nu}(d):=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\sqrt{2\nu}d\right)^{\nu}K_{\nu}(\sqrt{2\nu d}),$ (8)
where $K_{\nu}$ is a modified Bessel function of the second kind (Rasmussen and Williams, 2005). The number of basis coefficients used is 10% of the number of pixels, hence we have 90 basis coefficients for each region. For a less smooth input signal pattern, we can use a larger number of basis functions to increase the flexibility of the fitted functional coefficients.
All three Bayesian methods (BIOS, SBIOS0, SBIOSimp) use the same basis decomposition of the kernel function in (8), and all run for a total of 5000 MCMC iterations, where the last 1000 iterations are used for posterior inference. The prior for $\delta(s)$ is Bernoulli$(0.5)$. For the SGLD based methods (SBIOS0, SBIOSimp), we tune the step size $\tau_{t}$ across iterations using the formula $\tau_{t}=a(b+t)^{-\gamma}$ as in Welling and Teh (2011), and set $a=0.001,b=10,\gamma=0.55$ to ensure that through the 5000 MCMC iterations, the step size $s_{t}$ decreases from around $2\times 10^{-3}$ to $9\times 10^{-6}$. The initial values for all three Bayesian methods are identical, where $\delta(s)$, $\theta_{\beta,l}$ and $\theta_{\gamma,k,l}$ are set to 1, $\theta_{\eta,i,l}$ are set to 0, and all variance parameters $\sigma_{Y},\sigma_{\beta},\sigma_{\gamma},\sigma_{\eta}$ are set to 1.
4.1.1 Evaluation Criteria
To assess the variable selection accuracy, we compute the True Positive Rate (TPR) when the False Positive Rate (FPR) is controlled at $10\%$. Because of the selection variable $\delta(s)$, we can obtain a Posterior Inclusion Probability (PIP) from the $m=1,\dots,M$ posterior MCMC samples of $\delta(s)$,
$\text{PIP}(s)=\frac{1}{M}\sum_{m=1}^{M}\delta_{m}(s).$
Setting a threshold between 0 and 1 on $\text{PIP}(s)$ gives a mapping of the active pixels. Hence we choose 20 evenly-spaced points between 0 and 1 to fit an ROC curve with linear splines, and get the estimated TPR when FPR is at 10% from the fitted ROC curve. For MUA, we choose the cutoff on p-values to be the 20 quantiles of the Benjamini-Hochberg (BH) adjusted p-values corresponding to the probabilities at the 20 evenly-space points, and use the same method to obtain TPR when FPR controlled at 10%.
All three Bayesian methods (BIOS, SBIOS0, SBIOSimp) are implemented using Rcpp (Eddelbuettel and François, 2011) with RcppArmadillo (Eddelbuettel and Sanderson, 2014). Code for all simulations and all implementations of the proposed methods and competing methods can be found in the R package SBIOS. ( Github).
4.2 Simulation Results
We demonstrate the advantage of our proposed SBIOS algorithm concerning three objectives: variable selection accuracy, maximum memory usage, and scalability in terms of computing time as the sample size increases. The main results for selection accuracy, memory usage, and time scalability are shown in Figures 2, 4, and 5(a) respectively. We define the observed proportion (OP) at location $s_{j}$ to be $h(s_{j})=n^{-1}\sum_{i=1}^{n}\mathbf{1}_{\mathcal{V}_{i}}(s_{j})$, where $\mathcal{V}_{i}$ is the set of all observed locations for individual $i$.
4.2.1 Simulation I: Variable Selection
In Simulation I, to compare variable selection accuracy, when generating the individual masks, we provide results under two different missing data patterns as shown in Figure 1 (b) and (c). Figure 1 provides an illustration where in each pattern there exists a common area, and each location inside of this common area has observed data for all individuals. Outside of the common area, we allow missingness. This setting is motivated by the UK Biobank fMRI images where most missing voxels are at the brain periphery, and there is a large common area in the center of the brain with fully observed data. Outside of the commonly observed area, we allow the OP to be 0.5 or 0.1. Simulation I is performed under 8 scenarios: (i) OP of 0.5 or 0.1; (ii) $\sigma_{Y}=0.5$ or $\sigma_{Y}=1$; (iii) missing pattern follows Figure 1 (b) or (c). Changing the OP and missing patterns show the advantage of using imputation (SBIOSimp), and changing $\sigma_{Y}$ demonstrates algorithm performance under different signal-to-noise ratios.
Figure 2: True Positive Rate based on 100 replicated simulations when False Positive Rate is controlled at 0.1 when $n=3000,p=8100$. Missing Pattern I.
The selection accuracy result in Figure 2 shows that with Missing Pattern I, with the FPR controlled at 10%, SBIOSimp has much better TPR across different settings compared to the other methods. Within the 0-imputation based methods, the Bayesian methods (BIOS, SBIOS0) have similar performances, and are generally better than MUA, except for the case where $\sigma_{Y}=0.5$, OP is 0.5. In situations where estimating $\beta\delta$ becomes challenging, such as when the signal-to-noise ratio is low or the OP is low, spatially correlated methods are advantageous.
With Missing Pattern II, the three Bayesian methods outperform MUA in all settings. But since Missing Pattern II only generates missing pixels on non-active regions, there is smaller difference in FPR-controlled TPR between SBIOSimp and other methods, as shown in Section 2 of the Supplementary Material.
Comparing four sub-figures in Figure 2, we see that (i) as the OP decreases (comparing right subplots to left subplots), the TPR decreases for all methods, with MUA the most sensitive to the change in missing percentage; (ii) as the signal-to-noise ratio decreases ($\sigma_{Y}$ goes from 0.5 to 1, comparing top subplots to bottom subplots), the TPR decreases, again MUA being the most sensitive.
Figure 3: Use 95% cutoff on PIP for Bayesian active voxel selection.
Although Figure 2 provides a fair comparison of the power of each method given the same FPR tuned according to the true signal pattern, in practice, we prefer criteria for active voxel selection that can control the FDR well without knowing the true signal. When we use the 95% cutoff on the PIP for variable selection, the FDR and TPR are shown in Figure 3. This indicates that the 95% cutoff on PIP allows us to control FDR reasonably well below 0.05. The Supplementary Figure S4 provides a full comparison among three selection criteria: (i) Morris’ FDR control method (Meyer et al., 2015); (ii) MUA-based selection, i.e. choosing the cutoff on PIP such that the Bayesian method would have the same proportion of active voxels as identified by MUA BH-adjusted p-values; (iii) Directly applying the 95% cutoff on the PIP. The result in Figure S4 demonstrates that criteria (ii) and (iii) can both control the FDR well below 0.05, but the 95% threshold gives a better TPR for the Bayesian methods, hence we use this criterion as the cutoff on PIP in real data analysis.
4.2.2 Simulation II: Memory Usage
In simulation II, to compare memory usage, the generative setting sets $\sigma_{Y}=1$, missing pattern II, and an observed proportion of 0.5. The sample sizes are $n=3000,6000,9000,12000$, with fixed $p=3600$. SBIOSimp has similar memory usage to SBIOS0 except for a vector to save the imputed outcome, so we only compare MUA, BIOS, and SBIOS0.
To record maximum memory usage, we use the R function Rprof. In particular, we use summaryRprof with the option memory="tseries" where memory usage is recorded for segments of time. We take the maximum memory across all such segments as the maximum memory usage. The reported memory usage is for running each algorithm and does not include memory usage for reading data into memory. Reading memory is recorded separately from running memory. Detailed results can be found in Section 2 of the Supplementary Material. For MUA and BIOS, we directly read all data into memory first, and pass the address to BIOS. For SBIOS0 and SBIOSimp, each batch of data are read into R as file-backed matrices on the disk, and its address is passed to both SBIOS0 and SBIOSimp for sampling.
Figure 4: Memory cost in Mb with $n=3000,6000,9000,12000$, using 100 replications. The left panel shows maximum memory used, and the right panel shows total memory used. The mean, 97.5% and 2.5% quantile lines are also displayed.
Results are shown in Figure 4. The left panel in Figure 4 reflects the peak memory for each method, and is the main result of interest. In terms of maximum memory used over time, our proposed method SBIOS0 is comparable to MUA, which only performs linear regressions and is much more efficient than the standard Bayesian method, BIOS. In terms of total memory used over time (right panel in Figure 4), both Bayesian methods take much less total memory than MUA which needs to loop through $p$ locations in total, but our proposed method SBIOS0 still outperforms the other.
4.2.3 Simulation III: Time Scalability
In simulation III, we compare time scalability for BIOS, SBIOS0, and SBIOSimp, and exclude MUA since it only needs to run $p$ simple linear regressions and thus is linear in the number of voxels. For this simulation, the generative setting uses missing pattern II, and sample sizes $n=3000$ and $6000$, with a fixed number of voxels at $p=3600$.
(a) Computing time for Bayesian methods, in $\log_{10}(\text{seconds})$, for both $n=3000$ and $n=6000$ based on 100 simulations.
method SBIOS0 SBIOSimp
$n$ 3000 6000 3000 6000
OP=0.5, $\sigma_{Y}$ = 0.5 4.57 11.50 4.24 9.77
OP=0.1, $\sigma_{Y}$ = 0.5 4.58 7.78 5.73 9.56
OP=0.5, $\sigma_{Y}$ = 1 4.66 11.53 4.32 10.11
OP=0.1, $\sigma_{Y}$ = 1 4.61 7.72 5.76 11.44
(b) Ratio of speed for SBIOS0 and SBIOSimp when compared to BIOS, i.e.  time of BIOS over time of SBIOS0 (or SBIOSimp) under 4 cases and n = 3000 and 6000, based on 100 replicated simulation results.
Figure 5: Computational time comparison
As illustrated in Figure 5(a), the computing time (seconds) is reported on the log base 10 scale. Our proposed SBIOS0 and SBIOSimp are shown to have much better time scalability across all settings compared to BIOS which uses the full data and Gibbs sampling.
5 UK Biobank Application
In this real data application, we use 3D task fMRI data from the UK Biobank as the outcome and age as the single exposure variable. The three confounding variables are gender, age by gender interaction, and head size.
5.1 Data Preprocessing and Estimation Procedure
The outcome variable is the fMRI data obtained from an emotion recognition task. In this task, participants are required to identify which of two faces displaying fearful expressions (or shapes) presented at the bottom of the screen matches the face (or shape) displayed at the top of the screen (See details in the Hariri faces/shapes emotion task (Miller et al., 2016; Hariri et al., 2002), as implemented in the Human Connectome Project). The 3D fMRI data in MNI atlas space (Evans et al., 1994) is a rectangular prism of $91\times 109\times 91$ voxels, and we used a total of $n=38,639$ subjects with task fMRI. The NIFTI data is approximately 180 GB, and the processed RDS files are roughly 34 GB. The original NIFTI outcome data for one subject is around 3Mb, and the NIFTI mask data of binary format for one subject is around 26Kb. Using the R package RNifti (Clayden et al., 2023), the preprocessing time for one subject’s data, including directly loading the outcome and mask data from a DropBox folder, is around 2.3 seconds. Saving the preprocessed data into file-backed matrices is instantaneous. We use the Harvard-Oxford cortical and subcortical structural atlases (Desikan et al., 2006) for the analysis. After preprocessing, we have a total of 110 brain regions.
To define the analysis mask, we recall the definition of observed proportion (OP) at location $s_{j}$ to be $h(s_{j})=n^{-1}\sum_{i=1}^{n}\mathbf{1}_{\mathcal{V}_{i}}(s_{j})$, where $\mathcal{V}_{i}$ is the set of all observed locations for individual $i$. We define the group analysis mask as =$\mathcal{B}=\left\{s_{j}:h(s_{j})>0.5\right\}$, i.e., the area where each voxel has at least 50% observed data. As shown in Figure 6, the group analysis mask with completely observed data (purple area) covers significantly less area compared to $\mathcal{B}$, which has at least 0.5 observed proportion (blue area). In the complete observed data, large portions of the brain regions are missing, notably including the orbitofrontal cortex, the inferior temporal cortex, and the amygdala—regions crucial for emotion processing. In particular, the mask with complete observed data contains only 52 out of the 110 regions in the Harvard-Oxford atlas.
Figure 6: Analysis mask using an observed proportion threshold of 0.5 and an intersection mask (completely observed data). The purple area indicates 100% inclusion; the blue area indicates the mask with an observed proportion between 0.5 and 1.0.
After applying a common mask $\mathcal{B}$ with an observed proportion of 0.5, we end up with $\mathcal{B}$ that contains $p=121,865$ voxels. The image outcomes $Y_{i}(s)$ are standardized across subjects, and the continuous covariates are also standardized. For each region, we apply the Matérn kernel function but with different $\rho$ and $\nu$ parameters (8), to account for the different smoothness of each region. Both $\rho$ and $\nu$ are determined through grid search so that the empirical covariance of $Y(s_{j}),Y(s_{j}^{\prime})$ and the estimated covariance by the Matérn kernel have the smallest difference in Frobenius norm. The number of bases $L$ is chosen so that the cumulative summation $\sum_{l=1}^{L}\lambda_{l}$ accounts for 90% of the total summation of all eigenvalues, hence we have a total number of $L=16,879$. In Section 5.4, we provide a sensitivity analysis when the cutoff is based on 92% of the total summation.
The $n=38,639$ subjects are split into 50 batches of data, with each batch containing around $700$ to $800$ subjects. The subsample size is 200, and the step size decay parameters $a=0.0001,b=1,\gamma=0.35$ are chosen so that the step size roughly decreases from $7\times 10^{-5}$ to $5\times 10^{-6}$ over the 5000 MCMC iterations. We use the MUA results as the initial values for $\beta$ and $\gamma$ to run the SBIOS0 and SBIOSimp methods. The initial values for $\theta_{\eta,i,l}$ are set to 0 everywhere, and the initial values for $\delta(s)$ are set to 1 everywhere. The initial values for $\sigma_{Y},\sigma_{\eta},\sigma_{\beta},\sigma_{\gamma}$ are all set to 0.1. To check the convergence of SBIOSimp, we run three independent chains and use the Gelman and Rubin test. The $l_{2}$ norm of the residuals is computed for the last 1000 iterations from the three chains using 20% of the entire data. The point estimate of the Gelman-Rubin statistic is 1.03 with an upper confidence limit of 1.1, indicating that the MCMC chain has approximately converged.
5.2 Analysis Result
We present the top 10 regions identified by the SBIOSimp method in Table 1. This SBIOSimp result takes 20.5 hours to run on 1 core CPU HPC Cluster, and the max memory used is 14 GiB. Using the posterior sample for each region, we compute the Region Level Activation Rate (RLAR) as follows: denote $\mathcal{B}_{j}$ as the mask for region $j$, the RLAR is $\sum_{s\in\mathcal{B}_{j}}\delta(s)/|\mathcal{B}_{j}|$, where $|\mathcal{B}_{j}|$ is the total number of voxels in region $j$. Hence for each MCMC sample of $\delta(s)$, we obtain one sample of RLAR for all regions, and therefore obtain the posterior distribution of RLAR for all regions. We present histograms of the posterior distribution of the RLAR over the last 1000 MCMC iterations in Supplementary Material Section 3. Table 1 presents the top 10 regions with the highest posterior mean of RLAR and their 95% credible intervals. To present the marginal effect of each $\beta(s_{j})$, we also report the effect summing over each region computed separately for the positive and negative effects. Only voxels with a marginal posterior inclusion probability (PIP) over 95% are viewed as active voxels. Because the top 10 regions have no active voxels with positive effects, we only report the negative effect in Table 1. We also report the number of active voxels in Table 1. All active voxels in the top 10 regions have a negative effect, that is, as age increases, the cognitive ability to process tasks tends to decrease. This trend is also reflected in the raw data, as can be seen by the scatter plots of age against the average image intensity in Section 3 of the Supplementary Material. To interpret the numeric results in Table 1, take the Right temporal fusiform cortex, anterior division region as an example. On average, 99% of the voxels within this region have activity that is associated with age. Specifically, for a one standard deviation increase in age, there is a decrease of 14.81 in brain signal intensity, summed over all locations within this region.
Table 1: Top 10 regions ordered by Region Level Activation Rate (RLAR). The last columns reports the negative voxel effect size summed over each region i.e.  $\sum_{j\in\mathcal{B}_{j}}\mathbb{E}\left\{\beta(s_{j})|\beta(s_{j})<0\right\}$, and inside the bracket are the number of negative voxels. Only voxels with marginal inclusion probability greater than 0.95 are included, otherwise counted as zero effect voxel. Hence positive voxels are omitted due to low inclusion probability.
Region Name Size RLAR Neg Sum (Count)
Right intracalcarine cortex 634 1 -47.08 (632)
Right supracalcarine cortex 151 1 -8.38 (151)
Left temporal fusiform cortex, anterior division 301 1 -15.57 (299)
Left inferior frontal gyrus, pars triangularis 692 1 -48.86 (686)
Right inferior temporal gyrus, anterior division 314 1 -16.54 (303)
Left intracalcarine cortex 557 0.99 -40.29 (547)
Right temporal fusiform cortex, anterior division 276 0.99 -14.81 (263)
Left occipital pole 1977 0.99 -157.61 (1911)
left hippocampus 218 0.98 -12.51 (209)
Left inferior temporal gyrus, anterior division 346 0.98 -15.26 (316)
For a visual representation, Figure 7(a) shows the voxel level PIP in the sagittal plane. The highlighted red regions represent voxels with greater than 0.95 PIP. Figure 7(b) presents the effect size of $\beta(s)\delta(s)$, with the highlighted area in the range $(-0.06,-0.03)$. Note that from Figure 7(b), voxels with PIP greater than 0.95 also correspond to voxels with a larger absolute value of effect size. We notice that the active area (defined by voxel level PIP greater than 0.95) has a negative effect $\beta(s)$. This can also be validated by the scatter plot in Section 3 of the Supplementary Material, where the image outcome generally has a negative association with age across all individuals.
(a) Posterior inclusion probability (PIP). The color bar from black to red ranges in $[0.95,1]$. Sagittal plane. The first two sagittal slices are in the left hemisphere.
(b) Posterior mean of $\beta(s)\delta(s)$. The color bar from black to blue ranges from -0.03 to -0.06. The overlaying purple area is the mask of PIP greater than 0.95.
Figure 7: Illustration of results using a grayscale brain background image (ch2bet, Holmes et al. (1998)). Images are created using MRIcron (Rorden and Brett, 2000).
Based on our results, we have the following general interpretations: (i) when controlling for the confounders, age has a negative impact on the neural activity for emotion-related tasks; (ii) the negative effect reflected from each voxel is of very small scale, shown as in Figure 7(b), indicating a very low voxel level signal-to-noise ratio; (iii) the top 5 brain regions with the highest RLAR are (a) right intracalcarine cortex, right supracalcarine cortex, and left Temporal fusiform cortex, anterior division, all considered as critical areas for high-level visual processing including face recognition; (b) left temporal fusiform cortex, anterior division, a key structure for face perception, object recognition, and language processing (Weiner and Zilles, 2016); and (c) right inferior temporal gyrus, anterior division, an area for language and semantic memory processing, visual perception, and multimodal sensory integration (Onitsuka et al., 2004). These top 5 regions are also consistently identified in the sensitivity analysis when using half of the data as training data, see Section 5.4.
5.3 Comparing SBIOSimp with SBIOS0 and MUA
We compare the posterior mean of $\beta(s)$ given $\text{PIP}(s)\geq 0.95$ between SBIOS0 and SBIOSimp stratified by the observed proportions on these regions in Figure 8. Each point in Figure 8 represents the effect of age on the brain signal for one voxel in the brain. The 6 regions in Figure 8 are chosen with high missingness. Comparing blue dots (low observed proportion, $h(s_{j})\in[0.5,0.7)$), red dots (medium observed proportion, $h(s_{j})\in[0.7,0.9)$) with black dots (high observed proportion, $h(s_{j})\in[0.9,1]$), we can see that $\beta$ fitted with SBIOS0 on voxels with lower observed proportion tend to be closer to 0 or directly mapped to 0 according to $I(\text{PIP}(s)\geq 0.95)$, compared with SBIOSimp. This implies that by directly imputing missing outcomes with 0, SBIOS0 tends to put more shrinkage on the posterior mean of $\beta(s)$ compared to SBIOSimp. Hence SBIOS0 potentially has lower power than SBIOSimp to detect the signals, which could be justified by the simulation result shown in Figure 2.
A comparison of active $\beta$ selection between MUA and SBIOSimp is available in Table 2. For SBIOSimp, the active voxel selection is based on PIP greater than 95%. For MUA, for a fair comparison, the cutoff on BH-adjusted p-values is determined using the same proportion of active voxels selected by SBIOSimp, so that MUA and SBIOSimp select the same number of active voxels. Table 2 shows that MUA with 0-imputation tends to map more voxels of low observed proportion towards 0.
Table 2: Proportion of active voxels identified by MUA and SBIOSimp for each range of observed proportions (OP).
OP [0.5, 0.7) [0.7, 0.9) [0.9, 1]
MUA 0.15 0.39 0.62
SBIOSimp 0.27 0.46 0.62
Figure 8: Scatter plot of the posterior mean of $\beta(s_{j})I(\text{PIP($s_{j}$)}\geq 0.95)$ based on SBIOS0 (x-axis) and SBIOSimp (y-axis) on six selected regions with high missingness. Blue dots indicate voxels with observed proportion $h(s_{j})\in[0.5,0.7)$. Red dots indicate voxels with observed proportion $h(s_{j})\in[0.7,0.9)$. Black dots indicate voxels with observed proportion $h(s_{j})\in[0.9,1]$.
5.4 Sensitivity Analysis
To further validate the result reported in subsection 5.2, we conduct sensitivity analysis on different choices of the hyperparameter $IG(a,b)$ in the prior for $\sigma_{Y}^{2}$, the choice of $\sigma_{\beta}^{2}$, and the number of bases in the Gaussian kernel. The baseline setting for results in subsection 5.2 is $\sigma^{2}_{Y}\sim IG(0.1,0.1)$, hyperparameter $\sigma_{\beta}^{2}=0.01$, and the number of basis is based on 90% of total eigenvalues ($L=16,879$). In the sensitivity analysis, in case 1, the prior for $\sigma_{Y}^{2}$ is $IG(1,1)$, and $IG(0.1,0.1)$ for case 2. In case 3, we use $\sigma_{\beta}^{2}=1$. In case 4, we choose the number of bases based on 92% of total eigenvalues ($L=20,355$). For case 5, we vary the smoothness parameter $\nu$ in the Matérn kernel (8), and create a new kernel where $\nu$ in each region is set to 90% the value of $\nu$ in the standard kernel used in Section 5.2. Because Case 1 & 2 have very similar results, we report them together.
Table 3 presents the region-level difference between each sensitivity case and the final result presented in subsection 5.2. Cases 1 and 2 produce almost the same result and are reported together. We can see that the top four regions are consistently selected with the highest RLAR.
Table 3: Sensitivity Analysis of SBIOSimp on UK Biobank data. The RLAR for each region is reported. For the region names, (R) means right hemisphere, (L) means left hemisphere, AD means anterior division, PT means pars triangularis.
Region ID Region Name Cases 1&2 Case 3 Case 4 Case 5
24 (R) intracalcarine cortex 1.00 1.00 1.00 1.00
47 (R) supracalcarine cortex 1.00 1.00 1.00 1.00
85 (L) temporal fusiform cortex, AD 1.00 1.00 1.00 1.00
53 (L) inferior frontal gyrus, PT 1.00 1.00 1.00 1.00
14 (R) inferior temporal gyrus, AD 1.00 1.00 0.99 0.99
72 (L) intracalcarine cortex 0.99 0.99 0.99 0.99
37 (R) temporal fusiform cortex, AD 0.99 0.99 0.99 1.00
96 (L) occipital pole 0.99 0.99 0.99 0.99
105 (L) hippocampus 0.98 0.99 0.98 0.99
62 (L) inferior temporal gyrus, AD 0.98 0.98 0.99 0.99
6 Conclusion and Discussion
We propose a Bayesian hierarchical model for Image-on-Scalar regression, with a computationally efficient algorithm for posterior sampling, and apply our proposed method to the UK Biobank data. Our proposed model can capture high dimensional spatial correlation, account for the individual-level correlated noise, and provide uncertainty quantification on the active area selection through posterior inclusion probability. Our main computational contribution includes (i) employing a scalable Stochastic Gradient Langevin Dynamics (SGLD) algorithm for big data ISR, (ii) borrowing memory-mapping techniques to overcome memory limitations for large-scale imaging data, and (iii) utilizing individual-specific masks with an imputation approach to maximize the use of available imaging data.
Through extensive simulation studies, we compare different implementations of the proposed model, from the standard Gibbs sampling implementation (BIOS), the SGLD implementation with missing outcome imputed as 0 (SBIOS0), and the proposed imputation-based SGLD algorithm (SBIOSimp), with the standard Mass Univariate Analysis (MUA). The simulation results provide evidence that SBIOSimp can achieve better variable selection accuracy, with much smaller memory cost, and the computational time is scalable to large-scale data. In the UK Biobank application, we also validate our results by providing sensitivity analysis under different hyperparameters and kernel settings.
Although the performance of simulations and UK Biobank data indicates great potential in applying our method to other large-scale imaging applications, there are still several limitations. First of all, the underlying Gaussian process requires a user-defined kernel function. Currently, our data-adaptive way of choosing kernel parameters is still partially subject to the user’s discretion. Secondly, we assume that the individual effect $\eta$ shares the same kernel with $\beta$ and $\gamma$ mainly because of computational efficiency, but in practice there might be individual-level latent confounders that cannot be captured by estimating $\eta$. Another limitation comes from the SGLD algorithm. The main focus is to identify active areas and effect size of $\beta$, and SGLD can give us an accurate first-moment estimation of $\beta$. However, as the step size decreases to zero, the posterior MCMC sample may not truly recover the variance of $\beta$. A future direction of this work is to apply the proposed method to other large-scale imaging data such as the ABCD study Casey et al. (2018).
Appendix S1 Full Conditional Posterior Distributions
We provide the derivations of full conditional posterior distribution for each parameter. We will consider the case with completely observed data, and then provide the derivation for the case with imputing the missing values when considering individual masks.
We use $I_{d}$ to denote the identity matrix in $\mathbb{R}^{d\times d}$. We use $\mathrm{diag}\left\{D\right\}$ to denote a diagonal matrix whose diagonal vector is $D$.
S1.1 Posterior densities with completely observed data
Consider the basis expansion approximations of the GPs
$\beta(s)$ $\approx\sum_{l=1}^{L}\theta_{\beta,l}\psi_{l}(s)$ (9)
$\gamma_{k}(s)$ $\approx\sum_{l=1}^{L}\theta_{\gamma,k,l}\psi_{l}(s)$ (10)
$\eta_{i}(s)$ $\approx\sum_{l=1}^{L}\theta_{\eta,il}\psi_{l}(s)$ (11)
where $\theta_{\beta,l}\overset{ind}{\sim}N(0,\sigma_{\beta}^{2}\lambda_{l})$. Denote $D=\left(\lambda_{1},\dots,\lambda_{L}\right)^{T}$ as the vector of eigenvalues.
Let $Q\in\mathbb{R}^{p\times L}$ be the basis decomposition matrix, where $Q_{j,l}=\psi_{l}(s_{j})$. Here, to approximate the orthonormality of $\psi_{l}(s)$, $Q^{T}Q=I_{L}$, $Q$ is made an orthonormal matrix using QR decomposition.
Denote $Y^{*}_{i}=Q^{T}Y_{i}\in\mathbb{R}^{L}$ as the low-dimensional mapping of the $i$th image $Y_{i}\in\mathbb{R}^{p}$. After basis decomposition
$Y_{i}^{*}$ $=Q^{T}X_{i}\mathrm{diag}\{\delta\}Q\theta_{\beta}+\sum_{k=1}^{m}\theta_{\gamma,j}Z_{i,k}+\theta_{\eta,i}+\epsilon^{*}_{i}$ (12)
$\theta_{\beta}(s)$ $\sim N(0,\sigma_{\beta}^{2}\mathrm{diag}\left\{D\right\})$ (13)
$\theta_{\gamma,k}(s)$ $\sim N(0,\sigma_{\gamma}^{2}\mathrm{diag}\left\{D\right\})$ (14)
$\epsilon^{*}_{i}(s)$ $\sim N(0,I_{L})$ (15)
The posterior distribution can be derived as follows.
$\Sigma_{\beta(s)}|\text{rest}$ $=\left(\frac{1}{\sigma_{\beta}^{2}}\mathrm{diag}\left\{D\right\}^{-1}+\sum_{i=1}^{n}Q^{T}\mathrm{diag}\left\{X_{i}\delta\right\}Q\left(\frac{1}{\sigma_{Y}^{2}}I\right)Q^{T}\mathrm{diag}\left\{X_{i}\delta(s)\right\}Q\right)^{-1}$
$\theta_{\beta}(s)|\text{rest}$ $\sim N\left(\Sigma_{\theta_{\beta}(s)|\text{rest}}\left\{\frac{1}{\sigma_{Y}^{2}}\sum_{i=1}^{n}\left\{[Y_{i}^{*}-\theta_{\eta,i}-\sum_{k=1}^{m}\theta_{\gamma,k}Z_{i,k}]^{T}Q^{T}\mathrm{diag}\{X_{i}\delta(s)\}Q\right\}\right\},\Sigma_{\theta_{\beta}(s)|\text{rest}}\right)$
$\Sigma_{\gamma_{k}(s)}|\text{rest}$ $=\left(\frac{1}{\sigma_{\gamma}^{2}}D^{-1}+\sum_{i=1}^{n}\frac{1}{\sigma_{Y}^{2}}Z_{i,k}^{2}I_{L}\right)^{-1}$
$\theta_{\gamma,k}(s)|\text{rest}$ $\sim N\left(\Sigma_{\theta_{\gamma_{j}}(s)|\text{rest}}\left\{\frac{1}{\sigma_{Y}^{2}}\sum_{i=1}^{n}Z_{k,i}\left\{Y_{i}^{*}-\theta_{\eta,i}-\sum_{k^{\prime}\neq k}\theta_{\gamma,j}Z_{k^{\prime},i}-Q^{T}X_{i}\delta(s)Q\theta_{\beta}\right\}\right\},\Sigma_{\theta_{\gamma_{j}}(s)|\text{rest}}\right)$
$\delta(s)|\text{rest}$ $=1\times P_{1}(s)+0\times P_{0}(s)$
$P_{1}$ $\propto\exp\left\{-\frac{1}{2\sigma_{Y}^{2}}\left[\delta^{T}\mathrm{diag}\left\{\beta^{2}\|X_{i}\|_{2}^{2}\right\}\delta-2\left[\sum_{i}^{n}X_{i}(Y_{i}-\eta_{i}-\sum_{k=1}^{m}\gamma_{j}Z_{i,k})\right]^{T}\mathrm{diag}\left\{\beta\right\}\delta\right]\right\}$
Denote $D_{\delta}:=Q^{T}\mathrm{diag}\left\{\delta\right\}Q$. The log-Likelihood w.r.t. $\theta_{\beta}$ can be expressed as
$\log L(\theta_{\beta})$ $=-\frac{1}{2\sigma_{Y}^{2}}\sum_{i=1}^{n}\|Y_{i}^{*}-\theta_{\eta,i}^{*}(s)-X_{i}D_{\delta}\theta_{\beta}\|_{2}^{2}$
$=\frac{1}{\sigma_{Y}^{2}}\left(\sum_{i=1}^{n}(Y_{i}^{*}-\theta_{\eta,i}(s)-\sum_{k=1}^{m}\theta_{\gamma,k}Z_{i,k})X_{i}\right)^{T}D_{\delta}\theta_{\beta}-\frac{\sum_{i=1}^{n}X_{i}^{2}}{2\sigma_{Y}^{2}}\theta_{\beta}^{T}D_{\delta}^{2}\theta_{\beta}$ (16)
Based on the above derivations, we pre-compute the following summary statistics. These notations will help readers understand our code if one is interested.
  • XY_term_allsample: $\sum_{i=1}^{n}X_{i}Y_{i}(s)\in\mathbb{R}^{p}$
  • XqYstar_term_allsample: $Y^{*}_{L\times n}Z_{n\times m}\in\mathbb{R}^{L\times m}$
  • XXq_sumsq: $\sum_{i=1}^{n}X_{i}Z_{i}\in\mathbb{R}^{m}$
  • XcXq_sumsq: $\left(\sum_{i=1}^{n}Z_{i,j}Z_{i,[-k]}\right)_{j=1}^{q}\in\mathbb{R}^{(m-1)\times m}$
  • XY_eta_term: $\sum_{i=1}^{n}X_{i}Y_{i}(s)-\eta_{p\times n}X_{n\times 1}\in\mathbb{R}^{p\times 1}$
  • XqYstar_theta_eta_term: $Y^{*}_{L\times n}-\theta_{\eta,(L\times n)}\in\mathbb{R}^{L\times n}$
From the above derivation, we can see that the posterior covariance matrix for $\theta_{\beta}$ is a dense matrix, hence we only apply the SGLD algorithm on $\theta_{\beta}$.
S1.2 Imputing missing values using individual masks
In this subsection, we consider how to update the missing outcome $Y_{i}(s_{j})$ when considering the individual masks. Denote $Q_{i}$ to be the $i$-th rows in $Q$. For subject $i$ with missing voxel $s_{j}\notin\mathcal{V}_{i}$,
$Y_{i}(s_{j})$ $=X_{i}\delta(s_{j})Q_{j}\theta_{\beta}+\sum_{k=1}^{q}Q_{j}\theta_{\gamma,k}Z_{i,k}+Q_{j}\theta_{\eta,i}+\epsilon_{i}$ (17)
In practice, to avoid repeated access of the entire data set, we save all locations of missing voxels in a vector Y_imp. When accessing one batch of outcome data, the corresponding missing locations in $Y_{i}(s)$ is replaced by the imputed data saved in Y_imp.
Below is the detail algorithm for the imputation part. We use $\mathcal{V}_{i}^{c}$ to denote the collection of all missing voxels for subject $i$.
Algorithm 2 Update missing outcomes at one iteration for subjects in the $b$ batch.
1:Let Y_imp be the vector $\left\{Y_{i}(s_{j}):s_{j}\in\mathcal{V}_{i}^{c},i=1,\dots,n,j=1,\dots,p\right\}\in\mathbb{R}^{\sum_{i=1}^{n}|\mathcal{V}_{i}^{c}|}$. The length of Y_imp is the total number of missing outcomes across subjects and across locations.
2:Let $\mathcal{M}$ be a list of the index map that stores the location of missing voxels in Y_imp, i.e. the missing voxel $Y_{i}(s_{j})$ is located at the $k$-th element in Y_imp.
3:for  region $r=1,\dots R$do
4:     Extract index set $\mathcal{V}_{r}^{c}$, a list of index vectors, the $i$-th element is the vector $\mathcal{V}_{i}^{c}\cap\mathcal{B}_{r}$.
5:     Extract index set $\mathcal{M}_{r}$, a list of index vectors, the $i$-th element is the vector $\mathcal{V}_{i}^{c}\cap\mathcal{M}$.
6:     for subject $i$ with in batch $b$  do
7:         Extract index vector $\mathcal{V}_{r,i}^{c}$ which is the $i$th element in $\mathcal{V}_{r}^{c}$
8:         Extract index vector $\mathcal{M}_{r,i}$ which is the $i$th element in $\mathcal{M}_{r}$
9:         if  $\mathcal{V}_{r,i}^{c}$ and $\mathcal{M}_{r,i}$ are not empty set then
10:              Update Y_imp at locations $\mathcal{M}_{r,i}$ using (17) where $Y_{i}(s_{j}),s_{j}\in\mathcal{V}_{r,i}^{c}$.
11:         end if
12:     end for
13:end for
Note that when computing the likelihood, we need the imputed values from Y_imp. The reverse mapping from individual $i$ location $s_{j}$ to a particular $k$-th element in Y_imp is similar to the above algorithm.
Appendix S2 Additional Simulation Results
This section provides additional simulation results, including the TPR result for Missing Pattern II (Figure S1), the Morris FDR controlled selection results for Missing Pattern II (Figure S2), the reading memory cost (Figure S3), and the comparison among three active voxel selection criteria (Figure S4).
Below, Figure S1 provides the simulation results in Simulation I with missing Pattern II. The TPR is generally greater than with missing Pattern I since the missing pixels in missing Pattern II are all located outside of the active region. MUA presents the worst performance and SBIOSimp has the best performance across all settings, but the difference between SBIOSimp and the other competing methods is much smaller compared to the result in Simulation I with missing Pattern I.
Figure S1: True Positive Rate based on 100 replicated simulation results when False Positive Rate is controlled at 0.1 when $n=6000,p=8100$. Missing Pattern II.
Figure S2 provides the simulation results in Simulation I with missing Pattern II when applying Morris FDR control method to the Bayesian methods.
Figure S2: True Positive Rate and False Discovery Rate using Morris FDR control method, based on 100 replicated simulations, $n=3000,p=8100$. Missing Pattern II.
Figure S3 provides additional results on the reading memory in Simulation II: Memory Usage. Here, both SBIOS0 and SBIOSimp read $B$ batches of data in loops, and save them as file-backed matrices, hence the memory usage increase for reading memory is smaller compared to BIOS and MUA.
Figure S3: Maximum total reading memory (MB) for both $n=3000$ and $n=6000$ based on 100 replicated simulation results.
Figure S4 provides the comparison of FDR and TPR among three criteria: (i) Morris FDR control method (Morris, 2015); (ii) selection based on the MUA’s proportion of active voxels (determined by the BH-adjusted p-values less than 0.05); (iii) 95% cutoff on the PIP. Based on Figure S4, we can see that the Morris FDR control method cannot control the FDR below the target (0.05). The proportion of MUA can control the FDR well below 0.05 for the Bayesian methods, but the power is lower than the 95% cutoff on PIP.
(a) Morris’ FDR control method
(b) Use MUA’s proportion of active voxels to choose the cutoff on PIP.
(c) Use 95% cutoff on PIP.
Figure S4: True Positive Rates (TPR) and False Discovery Rates (FDR) using three criteria to choose the cutoff on PIP, based on 100 replicated simulations, $n=3000,p=8100$. Missing Pattern I.
Appendix S3 Additional UKBiobank Results and Figures
This section provides additional UKBiobank results. Figure S5 is a histogram of the RLAR posterior distribution for the top 10 identified regions. Figure S6 provides the scatter plot based on the observed data, indicating that there exists a negative association between age and brain signal intensity.
Figure S5: Posterior distribution of the region level active rate (RLAR) for the top 10 regions with highest mean region active rate.
Refer to caption Refer to caption
SBIOS0 SBIOSimp
est. Lower CI Higher CI p.value
-0.1731 -0.1938 -0.1523 $<0.001$
est. Lower CI Higher CI p.value
-0.1666 -0.1873 -0.1457 $<0.001$
Figure S6: Scatter plot of average image intensity(standardized) over the largest connected area against Age(standardized) for all individuals. The connected area is defined as follows: first we apply 0.7 cutoff on the inclusion probability to select active voxels with IP$>$0.7, then we search for the largest area with most connected active voxels. The scatter plot on the left is based on the inclusion probability obtained from SBIOS0 method, which happens to be a subset of the largest connected area obtained based on SBIOSimp. The right plot is based on GS-imputation.